OpenStudy (anonymous):

solve give exact solution. (7+√x)^2-14(7+√x)+22=0

OpenStudy (anonymous):

I got 27. is this correct?

OpenStudy (anonymous):

(7+√x)^2-14(7+√x)+22=0 \[49+14\sqrt{x}+x-98-14\sqrt{x}+22=0\] \[-27+x=0\] x=27 right @pleasecrystal

OpenStudy (anonymous):

thank you